slope of 5.1476x+5.1476y=15
Found 2 solutions by checkley77, drk:
Answer by checkley77(12844)

slope of 5.1476x+5.1476y=15
5.1476y=-5.1476x+15
y=-5.1476x/5.1476+15/5.1476
y=-x+2.914
Slope=-1 ans.

Answer by drk(1908)

The original question was to find the slope of:
5.1476x+5.1476y=15
step 1 - move the decimals 4 to the right to get
51476x+51476y=150000
step 2 - subtract 51476x from both sides to get
5.1476y = 150000 - 51476x
step 3 - divide by 51476 to get
y = -1x + 15000/51476
So, our slope is -1.
--
Alternate approach:
Ax + By = C
By = -Ax + C
y = -(A/B)x + C/B
Slope is opposite of X / y number.